Shift report 20000819 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sat Aug 19 2000 - 23:32:38 EDT

  • Next message: http account: "Shift report 20000820 00:00-08:00"

    ------------------------------------------------------------
    Date: Sat Aug 19 23:32:38 2000, Shift: 16:00-24:00
    Supervisor: JJG, with: EJK,KO
    Log book#4 pages 125 - 127
    
    Shift Summary:
    ------------------------------------------------------------
    Beam: Unsuccesfull attempts to inject. No beam at all. 
          Also RF problems.
    
    Magnets: D5 (4kG,700A);D1(410A);D2(410A)
    
    Specs: MRS at 0 deg, FS at 5 deg
    
    Malf: BB left :10 small tubes not working (HV mainframe?)
          T1 anode volt. cannot be raised. Mainframe? Access has
          been requested. When granted Ramiro should be called.
          3 H2 tubes bad.
    
    General: Need better record of angle and mag. settings. 
             EJK made a form to be filled by shift crew. KO will invent a window 
             in DAQ control so that info can be stored in RUN DB.
    
    Logbook: see estimate of events collected so far and guess 
             at number of p and p-bar.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sat Aug 19 2000 - 23:39:08 EDT